If the point (2,7) is on the straight line whose equation is ax+by=20, form the linear equation.
Answer by Alan3354(69443)

If the point (2,7) is on the straight line whose equation is ax+by=20, form the linear equation.
----------
Pick any value for a and you can colve for b.
--> an infinite # of lines.
eg,
a = 1
x + by = 20
2 + 7b = 20
b = 18/7
x + 18y/7 = 20
===============
a = 2
2x + by = 20
4 + 7b = 20
b = 16/7
2x + 16y/7 = 20
etc